# Pressure2Motion: Hierarchical Motion Synthesis from Ground Pressure with Text Guidance

### v1

@@ Abstract
We present Pressure2Motion, a novel motion capture algorithm that synthesizes human motion from a ground pressure sequence and text prompt. It eliminates the need for specialized lighting setups, cameras, or wearable devices, making it suitable for privacy-preserving, low-light, and low-cost motion capture scenarios. Such a task is severely ill-posed due to the indeterminate nature of the pressure signals to full-body motion. To address this issue, we introduce Pressure2Motion, a generative model that leverages pressure features as input and utilizes a text prompt as a high-level guiding constraint. Specifically, our model utilizes a dual-level feature extractor that accurately interprets pressure data, followed by a hierarchical diffusion model that discerns broad-scale movement trajectories and subtle posture adjustments. Both the physical cues gained from the pressure sequence and the semantic guidance derived from descriptive texts are leveraged to guide the motion generation with precision. To the best of our knowledge, Pressure2Motion is a pioneering work in leveraging both pressure data and linguistic priors for motion generation, and the established MPL benchmark is the first benchmark for this task. Experiments show our method generates high-fidelity, physically plausible motions, establishing a new state-of-the-art for this task. The codes and benchmarks will be publicly released upon publication.
Introduction
@@
Motion capture (MoCap) is a pivotal technology in digital animation production and robot control. However, classical motion capture systems are plagued by numerous limitations, such as reliance on wearable devices (in optical MoCap (kirk2004skeletal; guerra2005optical) and inertial MoCap (cloete2008benchmarking; vitali2020determining)), the high cost of equipment (particularly for optical motion capture), and dependence on illumination and visual information (in visual MoCap (mehta2020xnect)). These limitations have raised concerns regarding the cost, efficiency, and visual privacy protection of motion capture. To address this issue, we propose employing ground pressure mats in conjunction with simple textual prompts for motion capture, thereby eliminating the reliance on illumination and wearable devices.
@@
However, synthesizing human motion from ground pressure data is a highly challenging task due to its inherently underdetermined nature. Previous works on human pose prediction based on pressure data have only achieved success in scenarios involving large contact areas, such as predicting the poses of individuals lying in bed (clever2020bodies; wu2024seeing). Regressing human motions in standing poses, such as walking, from ground pressure data presents a far greater challenge. On the other hand, although significant progress has been made in the text-to-motion task (t2m; motiondiffuse), using text prompts as control signals is too unconstrained for a motion capture (MoCap) system. These methods rely on clean, manually specified, and kinematically abstract inputs, such as trajectories or keypoints (omnicontrol; zhong2025sketch2anim), to supply additional information beyond text input. Their architectures are tailored for these well-defined signals and are ill-equipped to handle a control modality that is noisy and physically grounded, such as ground pressure data.
@@
In this paper, we propose Pressure2Motion, a novel generative model that, for the first time, conditions motion synthesis on both high-level semantic text and low-level physical pressure as shown in Figure LABEL:fig:teaser. Our framework introduces a Dual-level Pressure Feature Extractor designed to interpret both the rich semantics of pressure (e.g., identifying which body part is in contact, its direction, and magnitude) and its temporal dynamics. This extractor decodes pressure sequences into two distinct control signals: an overall Pressure-Inferred Movement Trajectory and fine-grained Pressure-Induced Posture Shifts. These signals are then injected into a pre-trained motion diffusion model via a novel Hierarchical Pressure-Modulated Motion Synthesizer, which uses a ControlNet and a parallel Adapter module to ensure the final motion is both plausible and semantically consistent.
@@
To advance this novel research direction, we establish the MPL dataset. Constructed upon the extensive MotionPRO dataset (ren2025motionpro). MPL incorporates fresh, meticulously detailed textual annotations for each motion sequence, thereby establishing the inaugural large-scale, paired (text, pressure, motion) benchmark tailored specifically for this research task. We rigorously evaluate our proposed methodology on this dataset, and comprehensive experimental results demonstrate that Pressure2Motion is capable of generating human motions that are both highly realistic and physically coherent. This achievement heralds a new paradigm in non-visual, privacy-preserving motion capture.
@@
  • To the best of our knowledge, we are the first to formulate and address the novel task of generating full-body human motion from the highly sparse inputs of ground pressure and text prompts.
@@
  • By extending a pressure feature injection module on top of a pretrained text-to-motion model, both spatial alignment with ground pressure and semantic plausibility are achieved for the pressure-to-motion task.
@@
Related Work
Pressure-based Pose Estimation
@@
Traditional pose estimation methods rely on RGB cameras or inertial measurement units (IMUs), which lack reliable foot-ground contact cues and are unsuitable for privacy-sensitive scenarios like healthcare. Early pressure-based methods (yousefi2011bed; davoodnia2021bed; clever2020bodies) used single-frame pressure maps to estimate static postures, but were limited to predefined lying poses and low-resolution settings. Recent methods like PIMesh (wu2024seeing) extend this to short pressure sequences for mesh regression but struggle with dynamic motions like walking. Derived signals like center of pressure (COP)(funk2018learning; tripathi20233d) and foot contact(zhang2024mmvp) offer stability cues but are limited to quasi-static scenarios. Existing pressure datasets are mainly focused on in-bed or insole data, limiting coverage for free-moving activities. Most methods are constrained to short, static actions, with long dynamic sequences remaining a challenge.
@@
Multi-modal approaches, such as BodyPressure (clever2022bodypressure) and PressInPose (gao2024pressinpose), and others (yin2022multimodal; liu2022simultaneously; tandon2024bodymap) combine pressure with RGB, depth, or thermal sensors to improve pose accuracy, but still rely heavily on cameras, compromising privacy. MotionPRO (ren2025motionpro) dataset includes over 12.4M motion frames, significantly expanding coverage compared to earlier datasets like MoYo (tripathi20233d) and PSU-TMM100 (kissos2020beyond). They integrates plantar pressure with RGB-based motion capture, but still treats pressure as an auxiliary signal, underexploring its rich physical semantics. This limits its potential for contact-aware modeling and fails to leverage pressure sensing as a privacy-friendly alternative to vision-based systems.
Controllable Motion Synthesis
@@
Beyond generating human motion from natural language descriptions (guo2020action2motion; petrovich2021action; t2m; petrovich2022temos; jiang2023motiongpt; motiondiffuse; tevet2022human; Mofusion; chen2024text), recent research has increasingly explored controllable motion synthesis, where additional user-defined signals such as spatial constraints are introduced to steer or constrain the synthesis process. There are some approaches (posenetedit; ju2023humansd) that support partial pose constraints, while a prominent line of work focuses on controlling the position of key joints. PriorMDM (shafir2024human) finetunes MDM (tevet2022human) to allow control over end-effector trajectories, while GMD (karunratanakul2023guided) and Trace and Pace (rempe2023trace) guide motion through root joint trajectories. OmniControl (omnicontrol), MotionLCM (dai2024motionlcm) and MaskControl (maskcontrol) generalizes this idea, all leveraging ControlNet (zhang2023adding) for conditioning. Sketch2Anim (zhong2025sketch2anim) further extends this paradigm by leveraging both ControlNet and Adapter (rebuffi2017learning) to generate motion jointly conditioning on action semantics, keyposes and trajectories. While such methods provide flexible and expressive control, they rely on explicit, user-specified signals which are kinematic abstractions lacking physical grounding and cannot capture the continuous, frame-wise nature of real-world contact and balance phenomena.
@@
Figure 3: The Pressure2Motion pipeline. We first extract an overall Movement Trajectory $\mathbf{T}_{\text{traj}}$ and fine-grained Posture Shifts $\mathbf{S}_{\text{shift}}$ from pressure maps. These signals are then processed by our adapter branch to provide hierarchical control: a ControlNet encodes the trajectory for global guidance, while an Adapter Block fuses this with posture shifts for local refinement. The resulting features are injected into the pretrained MDM to generate plausible motion aligned with the pressure signals.
Dataset
To explore the role of pressure signals in motion synthesis, we introduce the MPL dataset, including Motion-Pressure-Language multimodal data for pressure-aware motion synthesis. This dataset is constructed by expanding upon the MotionPro dataset, a large-scale motion capture dataset that records ground pressure signals alongside full-body human motion. MPL dataset comprises motion sequences, ground pressure data, RGB video sequences, and corresponding textual descriptions, providing a rich resource for investigating the correspondence between ground pressure signals and human motion. Figure 2 shows samples from our dataset, including various motion sequences, their corresponding ground pressure data, and textual descriptions.
Dataset Statistics
@@
In the annotation phase, each RGB image sequence is labeled using Qwen2.5-VL (Qwen2.5-VL), a powerful vision-language model capable of processing and understanding long video sequences. For each motion sequence, we generate five textual descriptions with progressively decreasing levels of detail, resulting in a total of 104,720 descriptions across the dataset. By incorporating multiple descriptive levels, we enrich the variety of textual inputs, thereby enhancing our model’s generalization capability and enabling it to comprehend and synthesize a diverse array of motion styles and variations.
Motion Representation
Unlike MotionPro, which leverages the SMPL (SMPL) parametric human model, our approach employs a more comprehensive motion representation sourced from HumanML3D (t2m), an option that is more aptly tailored for motion synthesis frameworks. This representation includes pelvis velocity, local joint positions, joint velocities, joint rotations (in pelvis space), and binary foot contact indicators. This format ensures fine-grained control over the motion while maintaining compatibility with existing generative models.
@@
Method
Our objective is to synthesize human motion that aligns semantically with a given textual description while ensuring physical consistency with the contact dynamics encoded in the pressure distributions. Given a sequence of pressure maps $\mathbf{P}=\{\mathbf{p}^{i}\}_{i=1}^{N}$ over N frames, where each $\mathbf{p}^{i}\in\mathbb{R}^{H\times W}$ represents a frame-wise pressure map captured by ground pressure sensors, and a corresponding text description $c$, our goal is to synthesize a temporally coherent and physically realistic motion sequence $\mathbf{x}^{1:N}=\{\mathbf{x}^{i}\}_{i=1}^{N}$, where $\mathbf{x}^{i}\in\mathbb{R}^{D}$ represents a single pose and $D$ is the dimension of the representation, reflecting both semantic intent and pressure-based physical dynamics.
In this section, we first outline the basics of the Motion Diffusion Model (MDM), then detail two core components of Pressure2Motion: Dual-level Pressure Semantics Extraction and Hierarchical Pressure-Modulated Motion Synthesis. The complete pipeline is shown in Figure 3.
Preliminary: Motion Diffusion Model
Our model builds upon and expands the capabilities of the Motion Diffusion Model (MDM)  (tevet2022human), which was initially pretrained for generating motions from textual inputs, enabling its adaptation for pressure-sensitive motion synthesis. MDM is a denoising diffusion model that follows the DDPM framework (ho2020denoising). Given a motion sequence $\mathbf{x}$, it assumes $T$ noising steps modeled by the stochastic process:
@@
New motion sampling from MDM is performed iteratively, following the DDPM framework. At each time step $t$, the clean sample $\hat{\mathbf{x}}_{0}$ is predicted and then noised back to $\mathbf{x}_{t-1}$. This process is repeated from $t=T$ until the final clean motion, $\mathbf{x}_{0}$, is achieved. In our adaptation for pressure-aware motion synthesis, the pressure maps are integrated into the denoising process, providing crucial physical context that guides the generation of temporally consistent and physically realistic motions.
Dual-level Pressure Features Extraction
@@
The Pressure-Inferred Trajectory is essential for capturing the overall movement path and body alignment during motion. To extract this information, we utilize a feature extraction module $\mathcal{F}_{\text{traj}}$ following the architecture from MotionPro (ren2025motionpro), which includes a ResNet-based (he2016deep) pressure encoder, a temporal information processor, and a fully connected projection layer. Given the sparsity of pressure maps, where valid values are limited and primarily found under the feet during standing, the pressure encoder utilizes a compact ResNet architecture with small convolutional kernels to focus on the localized pressure regions, despite the large size of the pressure map. Temporal dynamics are captured using the temporal information processor, which combines a GRU (chung2014empirical) to model long-term dependencies with a self-attention mechanism to capture short-term correlations in the pressure sequence. The extracted spatial-temporal features are passed through a fully connected layer to generate a compact embedding $\mathbf{T}_{\text{traj}}$, which encapsulates the overall motion trajectory. The extraction of the Pressure-Inferred Trajectory is represented as the encoding of pressure information from each frame:
@@
where $\Delta\mathbf{P}$ represents the temporal difference in pressure maps and $\mathbf{e}$ is the grid-based positional encoding.
@@
Hierarchical Pressure-Modulated Motion Synthesis

### v2

# Pressure2Motion: Hierarchical Human Motion Reconstruction from Ground Pressure with Text Guidance

@@ Abstract
We present Pressure2Motion, a novel motion capture algorithm that reconstructs human motion from a ground pressure sequence and text prompt. At inference time, Pressure2Motion requires only a pressure mat, eliminating the need for specialized lighting setups, cameras, or wearable devices, making it suitable for privacy-preserving, low-light, and low-cost motion capture scenarios. Such a task is severely ill-posed due to the indeterminacy of pressure signals with respect to full-body motion. To address this issue, we introduce Pressure2Motion, a generative model that leverages pressure features as input and utilizes a text prompt as a high-level guiding constraint to resolve ambiguities. Specifically, our model adopts a dual-level feature extractor to accurately interpret pressure data, followed by a hierarchical diffusion model that discerns broad-scale movement trajectories and subtle posture adjustments. Both the physical cues gained from the pressure sequence and the semantic guidance derived from descriptive texts are leveraged to guide the motion estimation with precision. To the best of our knowledge, Pressure2Motion is a pioneering work in leveraging both pressure data and linguistic priors for motion reconstruction, and the established MPL benchmark is the first benchmark for this novel motion capture task. Experiments show that our method generates high-fidelity, physically plausible motions, establishing a new state of the art for this task. The codes and benchmarks will be publicly released upon publication.
Figure 1: By conditioning on pressure signals and text descriptions, Pressure2Motion reconstructs high-fidelity, physically realistic motions, addressing the challenge of synthesizing human motion from sparse and noisy pressure data.
1 Introduction
@@
Motion capture (MoCap) is a pivotal technology in digital animation production and robot control. However, classical motion capture systems are plagued by numerous limitations, such as reliance on wearable devices (in optical MoCap [25, 13] and inertial MoCap [6, 45]), the high cost of equipment (particularly for optical motion capture), and dependence on illumination and visual information (in visual MoCap [31]). These limitations have raised concerns regarding the cost, efficiency, and visual privacy protection of motion capture.
To address these limitations, we propose a novel MoCap paradigm: an algorithm that, at inference time, reconstructs full-body 3D motion using only a ground pressure mat in conjunction with simple textual prompts. Our goal is to create a system that can be deployed in privacy-sensitive environments without cameras or the need for users to wear invasive sensors. To learn the complex, ill-posed mapping from sparse pressure signals to full-body motion, our system is trained on a multimodal dataset containing synchronized pressure signals, ground-truth motion (captured via traditional sensors), and text. However, once trained, our system operates for inference without requiring any visual data or wearable devices.
@@
However, reconstructing human motion from ground pressure data is a highly challenging task due to its inherently underdetermined nature. Previous work on human pose prediction from pressure data has achieved success only in scenarios with large contact areas, such as predicting the poses of individuals lying in bed [4, 48]. Regressing human motions in standing poses, such as walking, from ground pressure data presents a far greater challenge. On the other hand, although significant progress has been made in the text-to-motion task [14, 15, 55, 43, 35, 22, 2, 16], using text prompts as control signals is too unconstrained for a MoCap system. More advanced controllable synthesis methods [54, 24, 10, 49, 56, 36, 32, 21] offer finer control but rely on clean, manually-specified, and kinematically abstract inputs (e.g., trajectories or keypoints) that have a direct geometric correspondence to the desired motion. Their architectures are ill-equipped to handle a physically grounded, noisy, and indirect control modality such as ground pressure, which lacks a simple kinematic mapping to full-body pose.
@@
In this paper, we propose Pressure2Motion, a novel generative model that, for the first time, conditions motion estimation on both high-level semantic text and low-level physical pressure as shown in Fig. 1. Our framework introduces a Dual-level Pressure Feature Extractor that interprets both the rich semantics of pressure (e.g., identifying which body part is in contact, its direction, and magnitude) and its temporal dynamics. This extractor decodes pressure sequences into two distinct control signals: an overall Pressure-Inferred Movement Trajectory and fine-grained Pressure-Induced Posture Shifts. These signals are then injected into a pre-trained motion diffusion model via a novel Hierarchical Pressure-Modulated Motion Synthesizer, which uses a ControlNet and a parallel Adapter module to ensure the final motion is both plausible and semantically consistent.
@@
To advance this novel research direction, we establish the MPL dataset. Constructed upon the extensive MotionPRO dataset [40]. MPL incorporates fresh, meticulously detailed textual annotations for each motion sequence, thereby establishing the inaugural large-scale, paired (text, pressure, motion) benchmark tailored specifically for this research task. We rigorously evaluate our proposed methodology on this dataset, and comprehensive experimental results demonstrate that Pressure2Motion can reconstruct human motions that are both highly realistic and physically coherent. This achievement heralds a new paradigm in non-visual, privacy-preserving motion capture.
@@
  • To the best of our knowledge, we are the first to formulate and address the novel task of full-body human motion capture from the highly sparse inputs of ground pressure and text prompts.
@@
  • We demonstrate how to effectively adapt a pretrained text-to-motion model for this novel MoCap task, achieving both spatial alignment with ground pressure and semantic plausibility.
@@
2 Related Work
2.1 Pressure-based Pose Estimation
@@
Traditional pose estimation methods rely on RGB cameras or inertial measurement units (IMUs)[30, 33, 46, 20, 23], which lack reliable foot-ground contact cues and are unsuitable for privacy-sensitive scenarios like healthcare. Early pressure-based methods [51, 9, 4] used single-frame pressure maps to estimate static postures, but were limited to predefined lying poses and low-resolution settings. Recent methods like PIMesh [48] extend this to short pressure sequences for mesh regression but struggle with dynamic motions like walking. Derived signals like center of pressure (COP)[11, 44] and foot contact[52] offer stability cues but are limited to quasi-static scenarios. Existing pressure datasets are mainly focused on in-bed or insole data, limiting coverage for free-moving activities. Most methods are constrained to short, static actions, with long dynamic sequences remaining a challenge.
@@
Multi-modal approaches, such as BodyPressure [5] and PressInPose [12], and others [50, 27, 42] combine pressure with RGB, depth, or thermal sensors to improve pose accuracy, but still rely heavily on cameras, compromising privacy. MotionPRO [40] dataset includes over 12.4M motion frames, significantly expanding coverage compared to earlier datasets like MoYo [44] and PSU-TMM100 [26]. They integrates plantar pressure with RGB-based motion capture, but still treats pressure as an auxiliary signal, underexploring its rich physical semantics. This limits its potential for contact-aware modeling and fails to leverage pressure sensing as a privacy-friendly alternative to vision-based systems.
2.2 Controllable Motion Synthesis
@@
Beyond generating human motion from natural language descriptions [14, 34, 15, 35, 22, 55, 43, 7, 2, 16], recent research has increasingly explored controllable motion synthesis, where additional user-defined signals such as spatial constraints are introduced to steer or constrain the synthesis process. There are some approaches [54] that support partial pose constraints, while a prominent line of work focuses on controlling the position of key joints. PriorMDM [41] finetunes MDM [43] to allow control over end-effector trajectories, while GMD [24] and Trace and Pace [39] guide motion through root joint trajectories. OmniControl [49], MotionLCM [8], MaskControl [36] and ACMDM [32] generalizes this idea, all leveraging ControlNet [53] for conditioning. Sketch2Anim [56] further extends this paradigm by leveraging both ControlNet and Adapter [38] to generate motion jointly conditioning on action semantics, keyposes and trajectories.  [21] takes input as sparse and flexible keyjoint signals to synthesize full-body motion. While such methods provide flexible and expressive control, they rely on explicit, user-specified signals which are kinematic abstractions lacking physical grounding and cannot capture the continuous, frame-wise nature of real-world contact and balance phenomena.
@@
Figure 3: The Pressure2Motion pipeline. We first extract an overall Movement Trajectory $\mathbf{T}_{\text{traj}}$ and fine-grained Posture Shifts $\mathbf{S}_{\text{shift}}$ from pressure maps. These signals are then processed by our adapter branch to provide hierarchical control: a ControlNet encodes the trajectory for global guidance, while an Adapter Block fuses this with posture shifts for local refinement. The resulting features are injected into the pretrained MDM to synthesize plausible motion aligned with the pressure signals.
3 Dataset
To explore the role of pressure signals in motion synthesis, we introduce the MPL dataset, including Motion-Pressure-Language multimodal data for pressure-aware motion synthesis. This dataset is constructed by expanding upon the MotionPro dataset, a large-scale motion capture dataset that records ground pressure signals alongside full-body human motion. MPL dataset comprises motion sequences, ground pressure data, RGB video sequences, and corresponding textual descriptions, providing a rich resource for investigating the correspondence between ground pressure signals and human motion. Fig. 2 shows samples from our dataset, including various motion sequences, their corresponding ground pressure data, and textual descriptions. More details can be found in the Sup. Mat.
3.1 Dataset Statistics
@@
In the annotation phase, each RGB image sequence is labeled using Qwen2.5-VL [1], a powerful vision-language model capable of processing and understanding long video sequences. For each motion sequence, we generate five textual descriptions with progressively decreasing levels of detail, resulting in a total of 104,720 descriptions across the dataset. By incorporating multiple descriptive levels, we enrich the variety of textual inputs, thereby enhancing our model’s generalization capability and enabling it to comprehend and synthesize a diverse array of motion styles and variations.
3.2 Motion Representation
Unlike MotionPro, which leverages the SMPL [28] parametric human model, our approach employs a more comprehensive motion representation sourced from HumanML3D [15], an option that is more aptly tailored for motion synthesis frameworks. This representation includes pelvis velocity, local joint positions, joint velocities, joint rotations (in pelvis space), and binary foot contact indicators. This format ensures fine-grained control over the motion while maintaining compatibility with existing generative models.
@@ 3.2 Motion Representation
4 Method
Our objective is to reconstruct human motion that aligns semantically with a given textual description while ensuring physical consistency with the contact dynamics encoded in the pressure distributions. Given a sequence of pressure maps $\mathbf{P}=\{\mathbf{p}^{i}\}_{i=1}^{N}$ over N frames, where each $\mathbf{p}^{i}\in\mathbb{R}^{H\times W}$ represents a frame-wise pressure map captured by ground pressure sensors, and a corresponding text description $c$, our goal is to synthesize a temporally coherent and physically realistic motion sequence $\mathbf{x}^{1:N}=\{\mathbf{x}^{i}\}_{i=1}^{N}$, where $\mathbf{x}^{i}\in\mathbb{R}^{D}$ represents a single pose and $D$ is the dimension of the representation, reflecting both semantic intent and pressure-based physical dynamics.
In this section, we first outline the basics of the Motion Diffusion Model (MDM), then detail two core components of Pressure2Motion: Dual-level Pressure Semantics Extraction and Hierarchical Pressure-Modulated Motion Synthesis. The complete pipeline is shown in Fig. 3.
4.1 Preliminary: Motion Diffusion Model
As established in Sec. 1, reconstructing full-body motion from sparse pressure is a severely ill-posed problem. Simple regression models (as shown in Tab. 1) often fail, producing unrealistic or physically implausible motions. We posit that this ill-posed reconstruction task is best solved using a generative prior. Diffusion models, such as the Motion Diffusion Model (MDM) [43], have proven to be exceptionally powerful priors for learning complex data distributions. Therefore, we adapt MDM not as a creative tool for synthesis, but as a powerful mechanism for reconstruction, using it to find the most physically plausible motion $\hat{\mathbf{x}}_{0}$ that matches the given pressure and text prompts.
Our model builds upon and expands the capabilities of the MDM, which was initially pretrained for synthesizing motions from textual inputs, enabling its adaptation for pressure-sensitive motion synthesis. MDM is a denoising diffusion model that follows the DDPM framework [19]. Given a motion sequence $\mathbf{x}$, it assumes $T$ noising steps modeled by the stochastic process:
@@ 4.1 Preliminary: Motion Diffusion Model
In our adaptation for pressure-aware motion synthesis, the pressure maps are integrated into the denoising process, providing crucial physical context that guides the synthesis of temporally consistent and physically realistic motions.
4.2 Dual-level Pressure Features Extraction
@@
The Pressure-Inferred Trajectory is essential for capturing the overall movement path and body alignment during motion. To extract this information, we utilize a feature extraction module $\mathcal{F}_{\text{traj}}$ following the architecture from MotionPro [40]. This module, detailed in the Sup. Mat., uses a ResNet [17] and GRU [3] to process the pressure sequence. The extracted spatial-temporal features are passed through a fully connected layer to generate a compact embedding $\mathbf{T}_{\text{traj}}$, which encapsulates the overall motion trajectory. The extraction of the Pressure-Inferred Trajectory is represented as the encoding of pressure information from each frame:
@@
where $\Delta\mathbf{P}$ represents the temporal difference in pressure maps and $\mathbf{e}$ is the grid-based positional encoding. This $\mathcal{F}_{\text{shift}}$ module is trained jointly, end-to-end, with the hierarchical pressure-modulated motion synthesizer described next.
@@
4.3 Hierarchical Pressure-Modulated Motion Synthesis